# Weak Existence and Uniqueness for Super-Brownian Motion with Irregular Drift

### v1

@@ Abstract
where $\dot{W}$ is space–time white noise and $h$ is a bounded drift with $h(0)\geq 0$. The proof relies on an extension of the duality relation of the super-Brownian motion, which allows us to treat a broad class of admissible drifts, including functions that are non-Lipschitz or discontinuous at zero.
@@ 1 Introduction
$d_{t}X_{t}=\frac{1}{2}\Delta X_{t}+h(X_{t})+\sqrt{X_{t}}\dot{W},\quad X_{0}=f,$ (1.1)
@@
for $b_{0}\geq\langle\nu_{1}-\nu_{2},1\rangle$, $b_{1}\in\mathbb{R}$ and finite measures $\nu_{1},\nu_{2}$ on $(0,\infty)$.
In this article, we establish weak well-posedness for solutions of (1.1) using duality methods. Hereby, we consider solutions in the space $\mathcal{C}(\mathbb{R}_{+},\mathcal{C}_{tem}^{+})$, which is the space of continuous functions taking values in
@@
We also allow random initial values that are distributed according to a measure in $\mathcal{P}_{p}(\mathcal{C}_{tem}^{+})$ (see Section 2.1). That means, we prove the following theorem.
@@
The main tool for the proof of Theorem 1.1 is the derivation of a duality relation (see Proposition 4.3). When $h(0)=0$, this relation also allows us to prove properties of the cozero set of the solution, even when $h$ has a discontinuity at zero.
Our assumptions on the drift function $h$ permit significant irregularities, including Hölder continuity with exponent less than $\frac{1}{2}$ and even discontinuities at zero. Several examples are provided in Section 5. Irregularities at zero are of particular interest, as this is precisely where the noise coefficient becomes non-Lipschitz. As a result, standard techniques such as the Dawson–Girsanov theorem can no longer be directly applied to establish weak uniqueness when the drift is less than $\frac{1}{2}$-Hölder continuous or exhibits discontinuities at zero. To the best of our knowledge, when the drift is discontinuous, neither weak uniqueness nor weak existence is known It turns out, that we can establish weak well-posedness of (1.1) for drift terms $h$ that lead to non-unique solutions or non-existence of solutions in the case of the one-dimensional SDE
$dx_{t}=h(x_{t})dt+\sqrt{x_{t}}dB_{t},\quad x_{0}=y,\quad t\geq 0,$
where $y\in\mathbb{R}$ and $(B_{t})_{t\geq 0}$ is a standard Brownian motion.
A similar behavior was also observed in [3] in the case of the Wright-Fisher SPDE.
Equation (1.1) with $h=0$ is the density of the super-Brownian motion in one dimension, which was extensively studied in the literature (see for instance [26, 15, 8, 10, 18, 23]). When the process is subject to some nonzero space dependent immigration, that is in the case where $h$ is continuous and does not depend on $X$ but instead on the spatial component $x$, pathwise non-uniqueness was established in [7].
When $h$ is continuous, (1.1) also fits in the setting of [25], where the existence of solutions is proven, but no uniqueness result was obtained for non-Lipschitz coefficients.
@@
Results on weak well-posedness for SPDEs with non-Lipschitz coefficients remain scarce. In the setting of Super-Brownian Motion and other measure valued branching processes a key method for establishing weak uniqueness is the duality approach. We will also employ this approach in our analysis. Since this method relies on guessing the correct duality functions it is sensitive to changes in the coefficients of the SPDE and thus one usually needs to assume a certain structure of the coefficients. The duality method was for instance used in [1, 21]. Hereby, the authors of [1] are using moment duality and in [21] the Laplace transform is used. In this article we will also use the Laplace transform as a starting point for our duality relation. Our approach is essentially an extension of the well known duality relation between the Super Brownian Motion and the log Laplace equation. More precisely, if $X$ denotes the solution to (1.1) with $h=0$ and $V(\mu)$ is the solution to
@@
for a finite measure $\mu$, then it is well-known (see for instance [8]) that
$\mathbb{E}[\exp(-\langle X_{t},\mu\rangle)]=\mathbb{E}[\exp(-\langle X_{0},V_{t}(\mu)\rangle)],\quad\forall t\geq 0,$
for all finite measures $\mu$. Furthermore, if $h\equiv a>0$, we get the following duality relation
@@
which was for example demonstrated in [19]. In this article, inspired by [21, 22] as well as [1], we are able to cover more complex immigration mechanisms $h$ by introducing a space-time jump process to (1.6). We outline this procedure in more detail in Section 4.
More recently, the authors of [3] extended the results of [1] for the one-dimensional Wright–Fisher SPDE by allowing certain irregular drift terms that were previously not covered. A key element in their analysis was the phenomenon of coming down from infinity in the dual branching–coalescing particle system, which is discussed in detail in [2]. Similarly, in this article, we employ the theory of initial trace solutions to the log-Laplace equation (see [6, 17, 20]) to extend the class of admissible drift functions $h$, allowing in particular for certain discontinuities at zero. The framework of initial trace solutions enables us to introduce jumps of infinite height in our dual process. In Section 4, we give some intuition on how this allows us to obtain a duality relation when $h$ has certain discontinuities at zero. We note that the theory of initial trace solutions is closely related to the phenomenon of coming down from infinity used in [3], both conceptually and technically.
In comparison to [3], we are able to control exponential moments of the number of jumps of the dual process in sufficiently small time intervals. This allows us to extend the class of admissible drift terms $h$ to include functions with $b_{0}>0$ and $b_{1}\leq 0$.
For this purpose, we introduce an artificial branching process that allows us to keep track of the number of jumps of the dual process. Again, this approach was inspired by [3].
@@
First, in Section 2, we introduce several function spaces and notation that is used throughout this article. In Section 3, we give a brief summary of results on certain non-linear parabolic equations, in particular the theory of initial traces and very singular solutions. Next, In Section 4, we outline the proof of Theorem 1.1. We then proceed to discuss several examples of admissible drift terms. In Section 6, we prove the duality relation (4.14). To this end, we introduce a branching process that allows us to control the moments of the number of jumps of the dual process in Section 6.1. We furthermore approximate the dual process by truncating the size of the jumps and prove the convergence of this approximation in Section 6.2. In Sections 7 and 8, we prove the weak existence and uniqueness of solutions by using the duality relation from Section 6. Finally, Section 9 is dedicated to the proof of Theorem 4.5, where we use the duality relation to prove properties of the cozero set of solutions to (1.1) under certain additional conditions.
@@ 2.1 Notation
    $\mathcal{P}(\mathcal{C}_{tem}^{+})$ denotes the set of all probability measures on $\mathcal{C}_{tem}$
@@ 3 Non-linear Parabolic Equations and Initial Trace
Furthermore, for $t<0$ we use the convention that $\mathcal{W}_{t}(A,\nu)=V_{t}(\mu)=0$ and $V_{0}(\mu)=\mu$.
@@ 3.1 Finite Measures
In [5] and [12, 22] the authors prove existence and uniqueness of solutions to (3.2) as well as several estimates. We summarized their results in this section. First, we get get existence and uniqueness by [22, Lemma 2.1], which we restated in the following proposition
@@
The following lemma is a well known fact about (3.2) (see for instance [22, Lemma 2.1]).
@@
The following lemma is [22, Lemma 2.6] and provides estimates that we shall use frequently.
@@
Finally, using [14, Lemma 2.3] we can infer the following lemma.
@@ 3.2 Initial Trace
In the next sections we shall rely heavily on the following result, which is taken from [20, Proposition 3.10] [17, Theorem 4] and [3, Section 2.1].
@@
    Let $((A_{n},\nu_{n}))_{n\geq 1}\subset\mathcal{T}$ such that $\eta^{A_{n},\nu_{n}}$ converges $m$-weakly to $\eta^{(A,\nu)}$, then it follows that $\mathcal{W}(A_{n},\nu_{n})$ converges to $\mathcal{W}(A,\nu)$ uniformly on compacts.
In particular, by Lemma 3.3 and Proposition 3.6 it follows that $V_{t}(n\delta_{x})y$ converges monotonically to $\mathcal{W}_{t}(x,0)(y)$ for all $t>0$ and $x,y\in\mathbb{R}$. Therefore,

### v2

@@ Abstract
where $\dot{W}$ denotes space–time white noise and $h$ is a bounded drift with $h(0)\geq 0$. The proof relies on an extension of the duality relation for super-Brownian motion, which allows us to treat a broad class of admissible drifts, including functions that are non-Lipschitz or discontinuous at zero. In particular, well-posedness is derived for certain drifts that are Hölder continuous at zero with exponent $\alpha\in(0,1)$. We also allow discontinuous drifts of the form $h(x)=b_{0}\mathbbm{1}_{x=0}+b_{1}\mathbbm{1}_{x>0},$ where $b_{0}\geq 0$, $b_{1}\in\mathbb{R}$. Additionally, if $h(0)=0$ and the initial condition is continuous and compactly supported, we show that the Lebesgue measure of the non-zero set of $X$ is finite. The proofs are based on duality. We use a log-Laplace equation which is perturbed by jump noise as the equation for the dual process and the jumps of the dual process are allowed to take infinite values. We believe that the results for the dual process are also of independent interest.
Under suitable assumptions on $h$ we also prove survival of $X$ with positive probability, using rescaling and comparison to the KPP-equation with branching noise.
@@ 1 Introduction
$d_{t}X_{t}(x)=\frac{1}{2}\Delta X_{t}(x)+h(X_{t}(x))+\sqrt{X_{t}(x)}\dot{W}(t,x),\quad X_{0}=f,x\in\mathbb{R},t\geq 0,$ (1.1)
@@
for $b_{0}\geq\langle\nu^{1}-\nu^{2},1\rangle$, $b_{1}\in\mathbb{R}$ and finite measures $\nu^{1},\nu^{2}$ on $(0,\infty)$.
In this article, using duality methods, we establish weak well-posedness for solutions of (1.1) for drifts satisfying conditions (1.21.4). A solution to (1.1) with the drift $h(x)=cx$, $c\in\mathbb{R}$, describes the density of the classical super-Brownian motion (SBM) with drift in one dimension, which has been extensively studied in the literature (see, for instance, [36, 21, 12, 14, 25, 32]), and for which weak well-posedness is well known. Our goal is to establish weak well-posedness for the much broader class of drifts satisfying conditions (1.21.4), beyond the classical case of $h(x)=cx$. We will comment further on this below.
We consider solutions in the space $\mathcal{C}(\mathbb{R}_{+},\mathcal{C}_{tem}^{+})$, which is the space of continuous functions taking values in
@@
We also allow random initial conditions distributed according to a measure in $\mathcal{P}_{p}(\mathcal{C}_{tem}^{+})$ (see Section 2.1). We now state the main results of the paper.
@@
A natural question concerning the properties of the solution is the following: How large can the cozero set of the solution be in the case where there is no upward drift at the zero points of the solution? That is, in the case when $h(0)=0$. In this setting, we prove that the cozero set of the solution to (1.1) has finite Lebesgue measure.
Theorem 1.2.
Suppose that almost surely, $X_{0}=f\in\mathcal{C}_{c}^{+}(\mathbb{R})$. Let $(X_{t})_{t\geq 0}$ be the solution to (1.1) with $h$ given by
$h(x)=\int_{0}^{\infty}(1-\operatorname{e}^{-\lambda x})\nu^{1}(d\lambda)+b_{1}\mathbbm{1}_{x>0},\quad x\in\mathbb{R},$ (1.6)
with $b_{1}\in\mathbb{R}$ such that $h(0)\geq 0$. Then there exists $T_{0}>0$ such that for all $t\in[0,T_{0}]$ we have
$\mathbb{P}(\mathrm{Leb}(\{x\in\mathbb{R}\colon X_{t}(x)>0\})<\infty)=1.$ (1.7)
Further fundamental questions concern the size of the closure of the cozero set, in particular whether the compact support property holds under the assumptions of Theorem 1.2. We conjecture the following.
Conjecture 1.3.
Suppose that almost surely, $X_{0}=f\in\mathcal{C}_{c}^{+}(\mathbb{R})$. Let $(X_{t})_{t\geq 0}$ be the solution to (1.1) with $h$ given by (1.6). Then $X_{t}$ is compactly supported with probability one for all $t\geq 0$.
Besides the cozero set and the support of the solution it is natural to investigate the longtime behavior, in particular the survival probability of the process.
Theorem 1.4.
Suppose that $h\neq 0$ is given by
$h(x)=\int_{0}^{\infty}(1-\operatorname{e}^{-\lambda x})\nu^{1}(d\lambda)+b_{1}\mathbbm{1}_{x>0},\quad x\in\mathbb{R},$
with $b_{1}\geq 0$ and let $(X_{t})_{t\geq 0}$ be the solution to (1.1) with $X_{0}=f\in\mathcal{C}_{c}^{+}(\mathbb{R})$ and $f\neq 0$. Then
$\mathbb{P}(\langle X_{t},1\rangle>0,\quad\forall t>0)>0.$
The main tool for the proof of Theorems 1.1 and 1.2 is the duality relation derived in Proposition 4.3.
Our assumptions about the structure of the drift function $h$ may seem peculiar, but they actually allow for significant irregularities of $h$ at zero, precisely the point where the noise coefficient $\sqrt{x}$ becomes non-Lipschitz and degenerate. For example, by choosing
$\nu^{1}=\nu^{2}\equiv 0,\quad b_{0}\neq b_{1},\quad b_{0}\geq 0,$ (1.8)
one obtains a drift $h(\cdot)$ that is discontinuous at zero. More precisely, this choice of parameters $\nu^{i},b_{i},i=1,2$, in (1.8), allows the drift $h$ to take any constant value at points where the solution $X$ of (1.1) is positive, and any constant non-negative value at points where $X$ equals zero. Furthermore, with $\nu^{2}\equiv 0$, for any $\alpha\in(0,1)$, by choosing an appropriate measure $\nu^{1}$ and suitable constants $b_{0},b_{1}$, one can obtain $h(x)$ such that it behaves like $cx^{\alpha}$ for small values of $x$, in the sense that
$\lim_{x\downarrow 0}\frac{h(x)}{x^{\alpha}}=c>0.$
To the best of our knowledge, weak uniqueness for such irregular drifts at zero has not been previously established. In fact, even the existence of a solution to equation (1.1) with a drift discontinuous at zero was not known. It is worth noting that standard techniques for handling drifts, such as the Dawson–Girsanov theorem, can no longer be directly applied to establish weak well-posedness when the drift has a Hölder exponent $\alpha<1/2$ at zero or exhibits a discontinuity at zero.
It is also noteworthy that we establish weak well-posedness of (1.1) for discontinuous drift terms $h$ that lead to non-uniqueness or non-existence of solutions for the one-dimensional counterpart of (1.1) given by
$dx_{t}=h(x_{t})dt+\sqrt{x_{t}}dB_{t},\quad x_{0}=0,\quad t\geq 0,$ (1.9)
where $(B_{t})_{t\geq 0}$ is a standard Brownian motion. To be more precise, we show that this equation admits non-unique solutions when $h(x)=\mathbbm{1}_{x>0}$ and no solution exists when $h(x)=\mathbbm{1}_{x=0}$. Notably, these are exactly the cases for which we prove weak well-posedness of (1.1) in this paper. The above examples in SPDE and SDE settings are discussed in more detail in Section 5.
The question addressed in this paper belongs to a broader area of well-posedness of SPDEs with irregular coefficients. In particular the SPDE
$d_{t}X_{t}(x)=\frac{1}{2}\Delta X_{t}(x)+h(X_{t}(x))+\sigma(X_{t}(x))\dot{W}(t,x),\quad X_{0}=f,\;x\in\mathbb{R},t\geq 0,$ (1.10)
where $\dot{W}$ is a space-time white noise, $h$ is a drift and $\sigma$ is a noise coefficient, has been extensively studied. Many results on well-posedness for this equation mainly dealt with the question:
How irregular can the drift be and what are conditions on the noise coefficient so that well-posedness for this equation holds?
This problem is also known under the name of regularization by noise, as it has been shown for some irregular $h$ that the equation without noise is ill-posed and when the noise is added, the equation becomes well-posed. When the noise is additive ($\sigma\equiv 1$), strong well-posedness was derived for some irregular function-valued drifts in [17], [18], and more recently even for some measure-valued and distribution-valued drifts in [1]. Later the result from [1] was extended in [11] to SPDEs driven by mutiplicative noises with $\sigma$ bounded away from zero (non-degenerate) and sufficiently regular. Weak well-posedness was established in [9] for (1.10) with additive noise for more irregular distributions than in [1]. The so-called path-by-path uniqueness for (1.10) with additive noise and with bounded drifts was derived in [8].
For SPDEs driven by multiplicative noise with non-Lipschitz and potentially degenerate noise coefficients, results remain pretty scarce. Strong well-posedness has been proved in [29] for an SPDE (1.10) where $\sigma$ belongs to the class of Hölder continuous functions with exponent greater than $3/4$ and the drift $h$ is Lipschitz continuous. As for the SPDE (1.10) with $\sigma$ being a Hölder function with exponent less than or equal to $3/4$ and zero drift, only weak uniqueness is known for very specific noise coefficients $\sigma$ (such as, for example, $\sigma(u)=u^{\gamma}$ with $\gamma\geq 1/2$, or $\sigma(u)=\sqrt{u(1-u)}$); see [30], [34]. In [34], weak uniqueness was established for the noise coefficient $\sigma(u)=\sqrt{u(1-u)}$ also in the presence of drifts of the form
$h(x)=c_{1}(1-x)-c_{2}x+c_{3}x(1-x),\quad x\in[0,1],\;{\rm for}\;c_{1},c_{2}\geq 0,c_{3}\in\mathbb{R}.$
Later, in [2], [27], [4], for $\sigma(u)=\sqrt{u(1-u)}$, the weak uniqueness results were extended to broader classes of drift terms. In [30], [34], [2], [4] the duality technique was used to prove uniqueness, whereas in [27] the Dawson-Girsanov theorem was applied.
Now we turn to the super-Brownian motion (SBM), corresponding to the setting $\sigma(X)=\sqrt{X}$, that is SPDE (1.1). As noted earlier, the uniqueness for drifts of the form $h(X)=cX$ has been known for a long time, and follows from the duality method. It is also worth noting, that when SBM is subject to nonzero space-dependent immigration, meaning $h$ is continuous and depends only on the spatial parameter $x$ and not on $X$, pathwise non-uniqueness was shown in [10].
Proving weak well-posedness for (1.1) requires establishing both weak existence and weak uniqueness. Weak existence, under the assumption that $h$ is continuous, follows from [35]. However, no uniqueness results were provided in [35] for non-Lipschitz coefficients.
@@
For SBM and other related measure-valued branching processes the duality approach is a widely used method for establishing weak uniqueness. We will also employ this approach in our analysis. Our approach is essentially based on an extension of the well-known duality relation between SBM and the corresponding log-Laplace equation. More precisely, let $X$ denote the solution to (1.1) with $h=0$ and let $V(\mu)$ be the solution to
@@
for a finite measure $\mu$, then it is well-known (see for instance [21, Lemma 1.2]) that
$\mathbb{E}[\exp(-\langle X_{t},\phi\rangle)]=\mathbb{E}[\exp(-\langle X_{0},V_{t}(\phi)\rangle)],\quad\forall t\geq 0,$
for all non-negative, continuous and compactly supported functions $\phi$. Furthermore, if $h\equiv a>0$, the duality relation becomes:
@@
which can be verified by using Itô’s formula. In this article, we extend this approach for more complex drift terms $h$ by introducing a space-time jump noise into the log-Laplace equation (1.11). The resulting duality relation is presented in Proposition 4.3. In fact, Proposition 4.3 is also used in the proof of Theorem 1.2.
We outline the construction of the dual process in detail in Section 4. We would like to highlight again the paper [4] which provided a number of very important insights that helped us in proving our main results. In [4] the well-posedness results of [2] were extended for (1.10) with $\sigma(u)=\sqrt{u(1-u)}$ for a large class of irregular drift terms, including some that are Hölder continuous and discontinuos at zero. The duality method, based on moment duality, was again used in [4], and the dual process was a branching–coalescing system of Brownian motions. A key element used in the analysis in [4] was the property of coming down from infinity of the dual branching–coalescing Brownian particle system. This property is discussed in detail in [3] in the case of for coalescing Brownian motions without branching. This phenomenon was crucial for establishing well-posedness in [4] for certain Hölder continuous and discontinuous at zero drifts. Analogously, in this article, we rely on the property of coming down from infinity, applied not to a particle system, but to solutions of the log-Laplace equation, whereas we use Laplace transform duality instead of moment duality. To successfully use this coming down from infinity property, we employ the theory of initial trace solutions to the log-Laplace equation (see [7, 24, 26]) which in turn allows us to extend the class of admissible drift functions $h$, particularly to those exhibiting discontinuities at zero. More precisely, this framework of initial trace solutions enables us to introduce jumps of infinite height in the dual process. In Section 4, we provide some intuition on how this leads to a duality relation even when $h$ has discontinuities at zero. Let us emphasize that the theory of initial trace solutions is conceptually closely related to the coming down from infinity property for particle systems used in [4].
Compared to [4], we are able to control exponential moments of the number of jumps in the dual process over sufficiently small time intervals. This control allows us to extend the class of admissible drift terms $h$ to include cases with $b_{0}>0$ and $b_{1}\leq 0$, whose counterparts were not covered in [4]. For instance, our results cover the drift $h(x)=b_{0}\mathbbm{1}_{x=0}$, that allows for upward drift only at points where solution equals to zero, an intriguing case for which no well-posedness results were previously known and which is not treated in [4].
We believe that establishing well-posedness for SBMs with irregular drifts, especially those discontinuous at zero, addresses an important open question in the theory of super-Brownian motions. Furthermore, we expect these results can be extended to more general classes of superprocesses, such as $\alpha$-stable superprocesses with branching mechanisms that are not necessarily quadratic. It would also be interesting to construct a branching particle system whose scaling limit yields a solution to equation (1.1) with a drift discontinuous at zero. The discontinuity of the drift at zero in (1.1) suggests that, in the pre-limiting particle system, there may be different subcritical or supercritical branching mechanisms or mass immigrations, depending on whether the spatial regions have non-negligible or negligible particle "density". These questions are left for future work.
@@
First, in Section 2, we introduce several function spaces and notation used throughout this article. In Section 3, we provide a brief summary of results on certain non-linear parabolic equations, focusing in particular on the theory of initial traces and very singular solutions. Next, in Section 4, we outline the proof of Theorem 1.1. Then, in Section 5, we discuss several examples of admissible drift terms. Section 6 is devoted to the construction of the dual process, its properties, and the duality relation. In Section 6.1, we introduce a branching process that in particular allows us to control the moments of the number of jumps in the dual process. We also approximate the dual process by truncating the size of the jumps and prove the convergence of this approximation in Section 6.2. In Section 6.3, we derive an approximate duality, and finally, Proposition 4.3 (the duality relation) is proved. In Sections 7 and 8, we use Proposition 4.3 and other results to prove the weak existence and uniqueness of solutions. Section 9 is dedicated to the proof of Theorem 1.2, where we again use the duality relation. Finally, in Section 10 we prove Theorem 1.4.
@@
Acknowledgments
We want to thank Zhenyao Sun and Ed Perkins for very helpful conversations. The work of the authors was supported in part by the ISF grant No. 1985/22.
@@ 2.1 Notation
    $\mathcal{P}(\mathcal{C}_{tem}^{+})$ denotes the set of all probability measures on $\mathcal{C}_{tem}$.
@@
    If $B\subseteq\mathbb{R}$ is an open set $\mathcal{M}_{Rad}^{+}(B)$ denote the set of non-negative Radon measures on $B$.
  12. 2.12.
@@ 3 Non-linear Parabolic Equations and Initial Trace
Furthermore, for $t<0$ we use the convention that $\mathcal{W}_{t}(A,\nu)=V_{t}(\mu)=0$ and $V_{0}(\mu)=\mu$, for $\mu\in\mathcal{M}_{F}^{+}$ and $\nu\in\mathcal{M}_{Rad}^{+}(A^{c})$.
@@ 3.1 Finite Measures
In [6] and [16, 31] the authors prove existence and uniqueness of solutions to (3.2) as well as several estimates. We summarized their results in this section. First, we get get existence and uniqueness by [31, Lemma 2.1], which we restated in the following proposition
@@
The following lemma is a well known fact about (3.2) (see for instance [31, Lemma 2.1]).
@@
The following lemma is [31, Lemma 2.6] and provides estimates that we shall use frequently.
@@
Finally, using [20, Lemma 2.3] we can infer the following lemma.
@@ 3.2 Initial Trace
In the next sections we shall rely heavily on the following result, which is taken from [26, Proposition 3.10], [24, Theorem 4] and [4, Section 2.1].
@@
[…]
In particular, by Lemma 3.3 and Proposition 3.6 it follows that $V_{t}(n\delta_{x})(y)$ converges monotonically to $\mathcal{W}_{t}(x,0)(y)$ for all $t>0$ and $x,y\in\mathbb{R}$. Therefore,